# Talking Wikidata: Communication patterns and their impact on community engagement in collaborative knowledge graphs

### v1

@@ Abstract
We study collaboration patterns of Wikidata, one of the world’s largest collaborative knowledge graph communities. Wikidata lacks long-term engagement with a small group of priceless members, $0.8\%$, to be responsible for $80\%$ of contributions. Therefore, it is essential to investigate their behavioural patterns and find ways to enhance their contributions and participation. Previous studies have highlighted the importance of discussions among contributors in understanding these patterns. To investigate this, we analyzed all the discussions on Wikidata and used a mixed methods approach, including statistical tests, network analysis, and text and graph embedding representations. Our research showed that the interactions between Wikidata editors form a small world network where the content of a post influences the continuity of conversations. We also found that the account age of Wikidata members and their conversations are significant factors in their long-term engagement with the project. Our findings can benefit the Wikidata community by helping them improve their practices to increase contributions and enhance long-term participation.
@@ 1 Introduction
Online peer-production communities are powerful tools with groups of self-organised people who work collaboratively towards a shared outcome. They offer innovation, efficiency, knowledge sharing, and democratisation of technology. Over the years, communities like those of Quora, [^1] https://www.quora.com/ Stack Overflow, [^2] https://stackoverflow.com/ Wikipedia, [^3] https://www.wikipedia.org/ and Wikidata, [^4] https://www.wikidata.org/wiki/Wikidata:Main_Page have made significant contributions to computer science. Collaborative knowledge graph (KG) communities, such as Wikidata, are particularly important as they provide structured knowledge from large datasets [55], and support various AI systems, including conversational agents and voice assistants [9]. The community plays a significant role in the maintenance, sustainability, and success of collaborative KGs [4]. However, these communities often face the challenge of long-term engagement from members [38, 35], with a small group responsible for the majority of contributions and decision-making [31, 34]. Therefore, it is essential to study the collaborative mechanisms behind contributors to understand what impacts their participation and contributions.
Studies related to collaborative KGs have primarily focused on emerging roles [31, 35] and lifespan [38] based on their editing activities and semi-structured interviews. However, a plethora of online discussion studies in peer-production showed that contributors’ discussions are a valuable source of insights [48, 15, 30]. Different peer-production systems, like an online encyclopedia, an open-source software system, and a KG, have different editing processes, which have raised concerns about the applicability of assumptions across different projects [10, 21, 43]. As a result, it is essential to investigate the different types of peer-production systems individually to better understand their unique characteristics and challenges.
To fill this gap, we investigated members’ behaviour in large-scale collaborative KGs. Specifically, we studied the Wikidata community and examined how members’ communication patterns affected the project. Wikidata is an open generic interest KG supporting numerous intelligent systems [49]. It has a diverse community of volunteers who collaborate online. To understand community interactions, we used a mixed-method approach with descriptive statistics, statistical tests, network analysis, as well as neural networks that combined graph and text embeddings. We analysed the behaviour of Wikidata contributors by answering:
@@
  • RQ2 - What impacts whether a discussion will receive a response?
  • RQ3 - Do discussions impact editors’ engagement?
@@
Our study indicates that discussion interactions between Wikidata editors form a network with a high clustering coefficient and low shortest path, suggesting a small world network. In this network, the conversation’s continuation is influenced by the network topology and the content of discussions. At the same time, the age of an editor in the community and their discussions have an influence on their engagement. Our findings can support Wikidata in improving its practices and tools. Additionally, our results can aid recommendation systems, such as those found in [53, 1], in providing improved suggestions for editor engagement.
@@ 2.2 The contributors
CkeckUsers These editors examine the IP address data of registered users, which is otherwise private information. Its primary purpose is to investigate sock-puppets.
@@ 2.4 The communication
Figure 1 shows an example of a Wikidata discussion. Both talk and communication pages follow a similar structure, with titles referred to as subjects separating the raised issues and a sequence of responses referred to as posts. Every post should end with the editors’ username or IP, in the case of anonymous contribution, and a timestamp. A set of posts under a subject title is referred to as a thread.
@@ 3.1 Online communities through the lens of network analysis
Previous studies on social media, online forums, and peer-production communities, investigated their communication patterns. Several studies focused on antisocial behaviour like anomaly detection (e.g., rumour, spamming, fraud, and cybercrime detection) [27, 36], sentiment detection (e.g., irony or sarcasm, and controversy detection) [18, 13], and manipulation (e.g. sock-puppets) [24, 41] to maintain healthy communities. In addition, other studies analysed the structure of communication networks. For example, before the advent of social networks ($2006$), instant messaging platforms like MSN engaged 245M members in 30B conversations. An analysis of the MSN members network [25], in which nodes represent members, and edges connect two members if they ever exchanged messages, found characteristics like a power law in degree distribution; high clustering coefficient, $0.1$; and one connected component including $>99\%$ of all nodes. These are similar to characteristics of a small world network [52], networks in which node distance grows proportionally to the logarithm of the number of nodes, rather than those of random networks [7].
@@
Similar research was conducted on a different type of community called Slashdot [14]. This is a news website where short news posts are published, and readers can comment and score them. The authors built a network in which nodes represented members and edges connected to members if they responded to each other. The Slashdot network showed that it had characteristics of a small world network, similar to the MSN network. They found a giant component containing $97\%$ of the nodes, clustering coefficient $0.05$ and average shortest path $4$. the study compared Slashdot with the characteristics of a typical social network. The authors explored the degree assortativity of the network, which showed whether highly connected users are preferentially linked to other highly connected ones or not [32]. Unlike social networks, which present strong assortativity with highly connected users connect with highly connected users, Slashdot presented neutral assortativity. This means that members do not show any preference in responding based on others’ ‘popularity’. Degree assortativity seems to be a fundamental difference between social and non-social networks [32] as shown from similar studies on Facebook and Twitter [11, 45].
@@
Prior work on the Wikipedia project, which shares similarities with the Wikidata project, found that it exhibits characteristics of a small world network [17]. In this analysis, nodes represented members and edges connected members who edited the same article. The network presented a clustering coefficient $0.9$ and average shortest path $5$ [17]. Furthermore, the authors observed neutral assortativity which is similar to the Slashdot community. However, another study built communication networks for the Wikipedia Wikiprojects, which are small groups where contributors collaborate to build articles under a specific topic like sport or biology, found that the networks are mostly disassortative with ‘popular’ editors communicating with less connected editors [37]. Finally, an investigation of the temporal effect of assortativity in social networks, including Wikipedia, revealed that the assortativity increased rapidly at the early stages of the network growth and then decayed to equilibrium after certain links were created [54].
@@
Previous studies explored the network structure in a variety of online communities. However, there is no prior work investigating the network dynamics of a collaborative KG. Many online communities exhibited characteristics of small world networks, which had significant implications for how information spread through the network. These implications include efficient information sharing, exposure to diverse ideas, enhanced problem solving, faster epidemic control (vandalism), and resilience to editor drop outs [17, 52]. Additionally, assortativity provided valuable indications of how people connect. A similar investigation of the Wikidata community could reveal valuable insight into information spread and collaboration within the community.
3.2 User characteristics in peer-production communities
@@
Previous studies explored behaviour patterns using user characteristics in peer-production communities to study their behavioural patterns [6, 20]. In the case of Wikipedia, prior work studied interactions between users when editing articles by building edit networks to connect editors working on the same articles. [6]. The authors used characteristics like the type of the edit (e.g., delete, undelete, restore) and text information (e.g., number of words added or deleted) to understand the work of individual authors and their positive and negative interactions. For the community behind Schema.org, an ontology engineering project, prior work explored discussion participation patterns [20]. The authors created user profiles using characteristics like the number of emails sent and replied to or the issues raised. The study revealed, that $10\%$ of users are responsible for $80\%$ of all contributions in community discussions. Additionally, the number of edits made was found to be positively correlated with the level of discussion activity [8]. This suggested that individuals who made many changes also participated in many discussions.
@@
Previous studies in Wikidata used editing activities and user characteristics to analyse editors’ behaviour. By considering the number of edits, type of edits, type of editors (registered, anonymous, or bot), and account age, Muller et al. [31] found that most editors made specialised contributions, and only a small piece contributed widely to the project. Similar results found by Piao et al. [34], who showed that $0.08\%$ editors contribute $80\%$ of all edits. Satasua et al. [38] focused on editors’ volume of edits over time and argued that editors with high account age and high volumes of edits had the highest impact on the project. In addition, Piscopo et al. [35] by combining editing activity with patterns in community discussions revealed that editors present diverse participation in editing and discussing.
@@
3.3 The factors that affect the behaviour of members in peer-production communities
Previous work explored how discussions influence peer-production communities. Studies in open source software communities [10] and Wikipedia [2, 3] showed that different types of conflicts in discussions, i.e. task, normative, process, etc. affected the projects differently. Normative conflicts (i.e. group function conflicts) negatively impacted the intention to remain in the project, and process conflicts (i.e. disagreeing on how to perform tasks) harmed contributors’ performance. Additionally, Arazy et al. [3] argued that often, task conflicts transition to process conflicts, affecting the quality of Wikipedia articles. Finally, Rychwalska et al. [37] argued that the quality of Wikipedia articles is related to the private communication between editors.
Prior work explored the factors influencing editors to drop out of Wikidata. Piao et al. [34] investigated engagement in Wikidata by studying the editing activities of editors. The authors used various statistical and pattern-based features like the number of edits, activity types, duration between edits, pair of entities editing and their characteristics. They found that the feature that had the greatest impact on predicting dropout was the total number of edits made in the last few months This is similar to findings in Wikipedia [12]. However, there is no prior evidence of how discussions influence editors’ behaviour in Wikidata.
@@ 4 Data
Wikidata provides monthly updated dumps of the KG in JSON and RDF format, along with the complete history of all edits in XML format. [^12] https://www.wikidata.org/wiki/Wikidata:Database_download For our work, we utilised the XML files that were updated until February 2023 with all activity logs and discussions.
@@ 4.1 Data preprocessing
We processed the raw text data to create our discussions dataset. Figure 2 presents the raw text for the example discussion shown in Figure 1. The figure displays that raw text follows a specific template for every page to identify the subjects of the threads(identified by $=$ subject $=$), and at the end of each post the username (identified by [[User: username|username]]), and the timestamp (formed as hour: minutes date month year (UTC)).
As in previous studies conducted in Wikipedia [48, 28], we phased several challenges during the process of processing the raw text discussions. We began by separating threads and posts on the page, and then extracted the information of the editors who wrote the posts, and when. Due to the large size of the Wikidata archived dumps, which amounted to $1.6$T, the minor changes in templates that occurred over the years, the differences in terms of discussion channels (e.g., some channels included multiple table templates, others changed their header/subject type over time) and human error, we expected a small percentage of false split in our dataset when separating threads and posts, and misspellings in the editors’ usernames.
We encountered three types of difficulties related to usernames: some posts missed the expected username signature at the end, likely due to editors inexperienced with Wikidata practices (this is similar to Wikipedia for $8.3\%$ of posts [28]); the multilingual nature of the projects made it challenging to follow the structure of the template in all languages, and track the username in raw text; and the freedom provided by Wikidata to create usernames with colours, highlights, and emojis, made it difficult to identify the usernames due to the ASCII encoding format related to this features we found in the raw text.
@@ 5 Methodology
We created a set of features for Wikidata editors similar to studies related to users’ behaviour in peer-production communities. In the case of Wikipedia [6], the authors investigated the number and type of user edits, and for Schema.org [20], the authors studied the activity level of users in the community. Furthermore, in the case of MSN and Slashdot, the authors studied netwrok metrics for their user networks [25, 14].
Table 3 shows the features and their descriptions. The history of edits dataset was used to measure each editor’s account age, which is the gap between their registered day and their last edit, and the number of edits. For the editors’ registered day we used a Wikimedia API. [^14] https://www.mediawiki.org/wiki/API:Main_page The discussions dataset was used to count the number of posts and classify editors into ‘asker’ and ‘responder’.
For the feature access level, we visited the Wikidata list of users page, [^15] https://www.wikidata.org/wiki/Special:ListUsers and searched each group of access level individually. We then assigned labels to our list of editors using the list of usernames for each access level.
Our feature list includes five centrality metrics, which are degree, eigenvector, closeness, betweenness and cluster coefficient. These metrics were measured based on the registered editors’ network described in detail below, Section 5.1.
Finally, we built a notion about who is considered an active or inactive editor for the label activity status. Wikidata classifies editors as active and inactive based on their monthly editing activities without considering editors permanently leaving or dropping out. In contrast, Sarasua et al. [38] highlighted that editors presented varying periods of absence with no edits, leading to a comeback or drop out. Therefore, the authors used edits’ percentiles to measure these gaps of inactivity and defined that after approximately $10$ months of no contributions, editors are considered permanently inactive. Similar to Sarasua et al. [38], we used edits’ percentile for the maximum gaps of inactivity of editors. Based on Figure 3, we chose the $85\%$ percentile, defining a threshold of $23$ months. We found that $67\%$ of editors were active and $33\%$ inactive.
During the ten years of Wikidata, we found several editors absent for $5$ to $9$ years but returned to contributing again (see Figure 3). The maximum gap of inactivity was $118$ months. These findings differ from the first four years of Wikidata, where Sarasua et al. [38] found the maximum gap of inactivity being $16$ months and $55\%$ of editors inactive.
Table 3: The lists of features used to investigate editors’ behaviour and their descriptions.
@@
talking type Label related to the way editors talk: ‘akser’ for editors mostly starting threads; ‘responder’ for editors mostly responding into threads.
@@
eigenvector It defines the centrality of a node as proportional to its neighbours’ importance. For example, for two nodes in a network with the same degree, the one connected to more nodes with a high eigenvector score should have a higher eigenvector score.
closeness It counts the average of shortest paths. Nodes with a high closeness score have the shortest distances to all other nodes.
betweenness It is based on the shortest path between nodes. Counts the number of times a node acts like a bridge.
clustering coefficient It quantifies how close its neighbours are to be a clique (complete graph).

### v2

@@ Abstract
We study collaboration patterns of Wikidata, one of the world’s largest open source collaborative knowledge graph (KG) communities. Collaborative KG communities, play a key role in structuring machine-readable knowledge to support AI systems like conversational agents. However, these communities face challenges related to long-term member engagement, as a small subset of contributors often is responsible for the majority of contributions and decision-making. While prior research has explored contributors’ roles and lifespans, discussions within collaborative KG communities remain understudied. To fill this gap, we investigated the behavioural patterns of contributors and factors affecting their communication and participation. We analysed all the discussions on Wikidata using a mixed methods approach, including statistical tests, network analysis, and text and graph embedding representations. Our findings reveal that the interactions between Wikidata editors form a small world network, resilient to dropouts and inclusive, where both the network topology and discussion content influence the continuity of conversations. Furthermore, the account age of Wikidata members and their conversations are significant factors in their long-term engagement with the project. Our observations and recommendations can benefit the Wikidata and semantic web communities, providing guidance on how to improve collaborative environments for sustainability, growth, and quality.
@@ 1 Introduction
Online peer-production communities are powerful tools with groups of self-organised people who work collaboratively towards a shared outcome. They offer innovation, efficiency, knowledge sharing, and democratisation of technology. Over the years, communities like those of Quora, [^1] https://www.quora.com/ Stack Overflow, [^2] https://stackoverflow.com/ Wikipedia, [^3] https://www.wikipedia.org/ and Wikidata, [^4] https://www.wikidata.org/wiki/Wikidata:Main_Page have made significant contributions to computer science. Collaborative knowledge graph (KG) communities, such as Wikidata, are particularly important as they provide structured knowledge from large datasets [55] in machine-readable way to support various AI systems, including conversational agents and voice assistants [9]. The community plays a significant role in the maintenance, sustainability, and success of collaborative KGs [4]. However, these communities often face the challenge of long-term engagement from members [38, 35], with a small group responsible for the majority of contributions and decision-making [31, 34]. Therefore, it is essential to study the collaborative mechanisms behind contributors to understand what impacts their participation and contributions.
A plethora of online discussion studies in peer-production communities showed that contributors’ discussions are a valuable source of insights [48, 15, 30]. At the same time, the variety of types and editing processes of peer-production systems, such as an online encyclopedia, an open-source software system, and a KG, have raised concerns about the applicability of assumptions across different projects [10, 21, 43]. As a result, it is essential to investigate the different types of peer-production systems individually to better understand their unique characteristics and challenges. For collaborative KGs, analysis of discussions is missing from the line of research, with previous work focusing primarily on emerging roles [31, 35] and lifespan [38] based on their editing activities and semi-structured interviews.
To fill this gap, we investigated members’ behaviour in large-scale collaborative KGs. Specifically, we studied the Wikidata community and examined how members’ communication patterns affected the project. Wikidata is an open generic interest KG supporting numerous intelligent systems [49]. It has a diverse community of volunteers who collaborate online. We analysed the behaviour of Wikidata contributors by answering:
@@
  • RQ2 - What factors affect whether a discussion will receive a response?
  • RQ3 - Do discussions affect editor engagement?
To understand community interactions, we used a mixed-method approach with descriptive statistics, statistical tests, network analysis, as well as neural networks that combined graph and text embeddings. Previous work in online communities, to understand the characteristics of contributors’ collaborations, analysed their communication using network analysis [25, 17, 11], and their behaviour patterns through members characteristics, such as volume of edits, posts, or account age [6, 20]. We followed similar practices to answer $RQ1$. Additionally, to answer $RQ2$ and $RQ3$ and identify the factors affecting communication and participation we develop a machine learning model combining features of members’ characteristics, their posts using text embedding vectors and their topology using graph embedding vectors.
@@
  • we recommend a list of practices to improve communication and participation
Our study indicates that discussion interactions between Wikidata editors form a network with a high clustering coefficient and low shortest path, suggesting a small world network. In this network, the network topology and the content of discussions influence the continuation of conversations. At the same time, the age of an editor in the community and their discussions influence their engagement. Our findings can support Wikidata in improving its practices and tools. Based on our findings, we recommend the development of new systems to track posts and assist in writing/asking about complex topics like KE. We also highlight the importance of improvements for assisting new members and enhancing engagement. Additionally, our results can aid recommendation systems, such as those found in [53, 1], in design, for example, by using advice from editors found to be long-lasting in the community and in disseminating new tools with the assistance of well connected and popular members.
@@ 2.2 The contributors
CheckUsers These editors examine the IP address data of registered users, which is otherwise private information. Its primary purpose is to investigate sock-puppets.
@@ 2.4 The communication
Figure 1 shows an example of a Wikidata discussion. Both the talk and communication pages follow a similar structure, with titles referred to as subjects separating the raised issues and a sequence of responses referred to as posts. Every post should end with the editors’ username or IP, in the case of anonymous contribution, and a timestamp. A set of posts under a subject title is referred to as a thread.
@@ 3.1 Online communities through the lens of network analysis
Previous studies on social media, online forums, and peer-production communities, has investigated their communication patterns. Several studies have focused on antisocial behaviour, such as anomaly detection (e.g., rumour, spamming, fraud, and cybercrime detection) [27, 36], sentiment detection (e.g., irony or sarcasm, and controversy detection) [18, 13], and manipulation (e.g. sock-puppets) [24, 41] to maintain healthy communities. In addition, other studies analysed the structure of communication networks. For example, before the advent of social networks ($2006$), instant messaging platforms like MSN engaged $245M$ members in $30B$ conversations. An analysis of the MSN members’ network [25], in which nodes represent members, and edges connect two members if they have ever exchanged messages, found characteristics like a power law in degree distribution; a high clustering coefficient, $0.1$; and one connected component including $>99\%$ of all nodes. These are similar to the characteristics of a small world network [52], in which node distance grows proportionally to the logarithm of the number of nodes, rather than those of random networks [7].
@@
Similar research was conducted examining the Slashdot site [14]. This is a news website for publishing. Readers can comment and score the posts. The authors built a network in which nodes represented members and edges connected to members if they responded to each other. The Slashdot network showed characteristics of a small world network, similar to the MSN network. The authors found a giant component containing $97\%$ of the nodes, a clustering coefficient of $0.05$ and an average shortest path of $4$ [14]. The study compared Slashdot with the characteristics of a typical social network using degree assortativity, which measured whether or not highly connected users were preferentially linked to other highly connected ones [32]. Unlike social networks, which present strong assortativity with highly connected users connect with other highly connected users, Slashdot presented neutral assortativity. This means that members did not show any preference in responding based on the ‘popularity’ of their peers. Degree assortativity seems to be a fundamental difference between social and non-social networks [32], as shown in similar studies on Facebook and Twitter [11, 45].
@@
Prior work on Wikipedia, a community which shares similarities with Wikidata, found that Wikipedia exhibits characteristics of a small world network [17]. In this analysis, nodes represented members and edges connected members who edited the same article. The network presented had a clustering coefficient of $0.9$ and an average shortest path of $5$ [17]. Furthermore, the authors observed neutral assortativity which is similar to the Slashdot community. However, another study built communication networks for the Wikipedia Wikiprojects, which are small groups where contributors collaborate to build articles under a specific topic, and found that the networks were mostly disassortative, with ‘popular’ editors communicating with less connected editors [37]. This may suggest that Wikiprojects have distinct characteristics from the other discussions, perhaps due to closely collaborating editors organising their work. This finding was similar to the distinction of topics discussed in Wikidata Wikiprojects [19] and the item and property talk pages, and project chat communication pages [22]. The topics in Wikiprojects are likely to be related to organisation and management, while the talk pages are related to KE activities. Additionally, an investigation of the temporal effect of assortativity in social networks, including Wikipedia, revealed that the assortativity increased rapidly at the early stages of the network growth and then decayed and remained stable after certain links were created [54].
@@
Previous studies have explored the network structure in a variety of online communities. However, there is no prior work investigating the network dynamics of a collaborative KG. Many online communities have been shown to exhibit characteristics of small world networks, which has significant implications for how information is spread through the network. This includes efficient information sharing, exposure to diverse ideas, enhanced problem solving, faster epidemic control (vandalism), and resilience to editor drop outs [17, 52]. Additionally, assortativity has provided valuable insights into how people connect. A similar investigation of the Wikidata community could reveal insight into information spread and group dynamics. This understanding could guide strategic decisions, improve engagement, enhance robustness (e.g., staying robust to random editor drop outs), and facilitate efficient information dissemination. For example, high assortativity is an indication of strong subgroups or cliques within the network. In this case, community managers might focus on engaging sub-communities separately. In contrast, low assortativity indicates an inclusive community where there is a flow of information between different levels of user engagement. In this case, broader community-wide engagement strategies could be more effective. These findings could support Wikidata in adopting better practices for effective communication and collaboration to produce high-quality outcomes and to encourage continued editor involvement.
3.2 Member characteristics in peer-production communities
@@
Previous studies have explored behaviour patterns using member characteristics in peer-production communities to understand their underlying mechanisms of collaboration [6, 20]. In the case of Wikipedia, prior work studied interactions between users when editing articles by building edit networks to connect editors working on the same articles. [6]. The authors used characteristics like the type of the edit (e.g., delete, undelete, restore) and content (e.g., number of words added or deleted) to understand the work of individual editors and their controversial interactions. For the community behind Schema.org, an ontology engineering project, prior work explored discussion participation patterns [20]. The authors created user profiles using characteristics like the number of emails sent and replied to or the issues raised. The study revealed, that $10\%$ of users are responsible for $80\%$ of all contributions in community discussions. Additionally, the number of edits made was found to be positively correlated with the level of discussion activity [8]. This suggested that individuals who made many changes also participated in many discussions.
@@
Previous studies in Wikidata used editing activities and user characteristics to analyse the behaviour of editors. By considering the number of edits, type of edits, type of editor (registered, anonymous, or bot), and account age, Muller et al. [31] found that most editors made specialised contributions, and only a small piece contributed widely to the project. Similar results were found by Piao et al. [34], who showed that $0.08\%$ editors contributed $80\%$ of all edits. Sarasua et al. [38] focused on the volume of edits for over time and argued that editors with high account age and high volumes of edits had the greatest impact on the project. In addition, Piscopo et al. [35] by combining editing activity with patterns in community discussions, revealed that editors presented diverse participation in editing and discussing. Our analysis combines participation levels with the duration of engagement, the editing rights of editors (Table 1) and discussion patterns to provide more insights into the behaviour of editors and engagement patterns. This understanding can be used to tailor engagement strategies
@@
3.3 The factors affecting members in peer-production systems
Previous work has explored how discussions influence peer-production communities. Studies in open source software communities [10] and Wikipedia [2, 3] have shown that different types of conflicts in discussions, such as task, normative, and process conflicts, affected the projects differently. Normative conflicts (i.e., group function conflicts) negatively impacted the intention to remain in the project, and process conflicts (i.e., disagreeing on how to perform tasks) harmed contributors’ performance. Additionally, Arazy et al. [3] argued that often, task conflicts transition to process conflicts, affecting the quality of Wikipedia articles. Finally, Rychwalska et al. [37] showed that the quality of Wikipedia articles was related to the private communication between editors.
Prior work has also explored the factors influencing editors to drop out of Wikidata. Piao et al. [34] investigated engagement in Wikidata by studying the editing activities. The authors used various statistical and pattern-based features, such as the number of edits, activity types, duration between edits, pair of entities editing and their characteristics. They found that the feature that had the greatest impact on predicting dropout was the total number of edits made in the last few months. This is similar to findings in Wikipedia [12]. However, there is no prior evidence of how discussions influence editors’ behaviour in Wikidata.
@@ 4 Data
The Wikimedia Foundation supports publicly available dumps, updated monthly, for all its hosted projects. In our work, we used the XML dumps with the history of all edits. [^12] https://www.wikidata.org/wiki/Wikidata:Database_download For our work, we utilised the XML files that were updated until February 2023 with all activity logs and discussions.
@@ 4.1 Data preprocessing
During the data preparation, we processed raw text data. Figure 2 shows an example of raw text from a discussion page and Figure 1 the respective front page for the same discussion. In the figure, we can see that raw text follows a template for every page using specific punctuation to identify the subjects of the threads(i.e., $=$ subject $=$), and the username (i.e., [[User: username|username]]) and timestamp (i.e., hour: minutes date month year (UTC)) at the end of each post. Similar to previous studies in Wikipedia [48, 28], we phased multiple challenges in the process of the raw text discussions. We processed the text taking place on a discussion page, starting with separating threads and posts on the page, and thereafter extracting the information on the contributors who wrote the posts, and when. Due to the large size of the Wikidata archived dumps, $1.6T$, the minor changes in templates that happened in the project over the years, the differences in terms of discussion channel (e.g., some channels include multiple table templates, others change their header/subject type over time), and human error, we expected a small percentage of error in our dataset when separating threads and posts, and misspellings in the editors’ usernames.
We encountered three types of difficulty related to usernames: some posts missed the expected username signature at the end, probably due to editors who were inexperienced with Wikidata practices (this is similar to Wikipedia for $8.3\%$ of posts [28]); the multilingual nature of the projects made it challenging to follow the structure of the template in all languages and track the username in raw text; and the freedom provided by Wikidata to create usernames with colours, highlights, and emojis made it challenging to identify the usernames due to the ASCII encoding format related to this features that we found in the raw text.
@@ 5 Methodology
We created a set of features for Wikidata editors similar to studies related to the behaviour of contributors in peer-production communities. Previous research on Wikidata examined the duration of editors’ participation [38]. We introduced the feature account age to analyse the various membership lengths in months that editors have within this community, and the activity status to distinguish the active editors from those dropped out. Furthermore, in Wikidata and Wikipedia, previous studies analysed the volume of edits for the editors [6, kittur2007he, 38]. Similarly, we included the features number of edits, and number of posts. In ontology engineering studies, such as Schema.org [20], the authors analysed the characteristics of editors. In this vein, we introduced the talking type, which determines whether an editor mostly starts a discussion or responds to others, and the access level, which specifies the role of editors in the community. Finally, in the case of MSN and Slashdot [25, 14], the authors studied network metrics for their user networks, which led us to study centrality metrics, such as degree, eigenvector, closeness, betweenness, and clustering coefficient.
Table 3 shows the features and their descriptions. The history of edits dataset was used to measure each editor’s account age, which is the gap between the day they registered and their last edit, and the number of edits. For registered days, we used a Wikimedia API. [^14] https://www.mediawiki.org/wiki/API:Main_page The discussions dataset was used to count the number of posts and classify editors into ‘asker’ and ‘responder’.
Table 3: The lists of features used to investigate the behaviour of editors and their descriptions.
@@
talking type Label related to the way editors talk: ‘asker’ for editors mostly starting threads; ‘responder’ for editors mostly responding to threads.
@@
eigenvector Defines the centrality of a node as proportional to its neighbours’ importance. For example, for two nodes in a network with the same degree, the one connected to more nodes with a high eigenvector score should have a higher eigenvector score.
closeness Counts the average of shortest paths. Nodes with a high closeness score have the shortest distances to all other nodes.
betweenness Based on the shortest path between nodes. Counts the number of times a node acts like a bridge.
clustering coefficient Quantifies how close its neighbours are to be a clique (complete graph).
@@
For the feature access level, we visited the Wikidata list of users page, [^15] https://www.wikidata.org/wiki/Special:ListUsers and searched each access level group individually. We then assigned labels to our list of editors using the list of usernames for each access level.
Our feature list includes five centrality metrics, which are degree, eigenvector, closeness, betweenness and cluster coefficient. These metrics were measured based on the registered editors’ network, described in detail below, Section 5.1.
Finally, we built a notion about who was considered an active or inactive editor for the label activity status. Wikidata classifies editors as active and inactive based on their monthly editing activities without considering editors permanently leaving or dropping out. In contrast, Sarasua et al. [38] highlighted that editors presented varying periods of absence with no edits, leading to a comeback or drop out. Therefore, the authors used the percentiles of editors for gaps of inactivity and defined that after approximately $10$ months of no contributions, editors were considered permanently inactive. Similar to Sarasua et al. [38], we used the percentile of editors for the maximum gaps of inactivity of editors. Based on Figure 3, we can see that the majority of editors had a maximum gap of inactivity of less than $20$ months. We chose the $85\%$ percentile, defining a threshold of $23$ months. According this definition, we found that $67\%$ of editors were active and $33\%$ inactive.
@@
During the ten years of Wikidata, we found several editors who were absent for five to nine years but returned to contributing (see Figure 3). The maximum gap of inactivity was $118$ months. These findings differ from the first four years of Wikidata, where Sarasua et al. [38] found the maximum gap of inactivity being $16$ months and $55\%$ of editors inactive. There are numerous reasons, from personal issues to global events, why editors present these long absences. Further analysis could reveal insights into engagement patterns. It is interesting to note that editors with long periods of inactivity present a variety in the number of edits. Figure 3 shows that editors with almost seven years of absence can have from $10$ to $100K$ edits.